11. The owner of a ski apparel store in Winter Park, Colorado, must make a decision in July regarding the number of ski jackets to order for the following ski season. Each ski jacket costs $54 each and can be sold during the ski season for $145. Any unsold jackets at the end of the season are sold for $45. The demand for jackets is expected with a average rate of 80 The store owner can order jackets in lot sizes of 10 units a. How many jackets should the store owner order if she wants to maximize her b. What are the best-case and worst-case outcomes the owner may face on this c. How likely is it that the store owner will make at least $7,000 if she implements d. How likely is it that the store owner will make between $6,000 to $7,000 if she expected profit? product if she implements your suggestion? your suggestion? implements your suggestion?

If could be done using excel or google sheets that would be appreciated.
Answer #1

This problem can be solved by using MS Excel Solver as shown below:

Demand Lot Probaty Conditional loss Ex 1 Polsson averag 2 profit per Jacket 3 loss per unsold Jacket İholla: if the Qty demanded is same as stocked then the loss Is zero Each unsold usinit makes a loss of $9 (54-45) If the demand is more than what is stocked, then each last opportunity leads to a loss ol 45 6.57172E-06 46 1.14291E-05 47 し94537E-05 4D 3.24229E-05 49 5.29354E-05 50 .45966E-05 51 0.000132857 405 0.002661546 396 0.004525914 0.0075286 5 Solver Stock 6 Real Stock 7 Expected loss Author: Since the stodong has to be n the multples of 10, se stock wil be constrained as integer in the sover and the real stock is solver stock 78 0.012255061 69 0.019533151 360 0.030490772 51 0.045632946 342 0.05990343 333 0.102737911 324 0.148090682 315 0.209421167 306 0.290625293 297 0.395898232 288 0.529519056 279 0.055554692 270 0,89148992 261 1.137801435 252 1.417505681 243 1.735721242 234 2.089294088 225 2.472537382 216 2.877134408 207 3.292243352 198 3.704826279 189 4.100202997 10 4.462805983 171 4.777098094 162 5.023513733 15 5.204550034 144 5.295567603 135 5.295567603 0.03040772 Salver Parameters 857 jholla: Expected loss is the sun of all the expected losses from the right l side cxolumns. 54 0.00045707 We want to minimize this loss By Changrg Varisble Cels 56 0.000949756 57 0.001332991 58 0.001838608 59 0.002453028 00 0,003324037 Subjact to the Constrants $8$5-integer 62 0.005625023 63 0.00T112836 64 0.008928607 65 0.010989055 66 0.013320057 67 0.015904557 Reset Al 69 0.021694196 70 0.024793357 71 0.027936138 72 0.031040209 5eect a Solving Method GRG Nonlner oing Method Selact the GRG Noninear engne for Solver Problems that are smooth roninear. Select the LP Smplex engine for inear Solver Problems, and select the Evolutionary engine for Solver problems that are 74 0.036774775 75 0.039226427 Sheet1 Ready Sove

The excel sheet with the formula is shown below:

Demand Lc Probability Conditional loss Expected loss profit per jacket loss per unsold jacket 2 145-54 3 6 7 solver Stock Reaa) From the above workings, we find that the optimum number of jackets to store is 90

b) The best case outcome when stocking 90 is when the demand is 90. The total profit earned is 90*(145-54)= $8190

The worst case can be as high as any demand as we are considering the opportunity cost which is the opportunity of making a profit of $91 by selling a jacket

c) The profit for each jacket sold is $91 and loss due to an unsold jacket is $9.

Profit when the demand is 78 and stock is 90 = 78*91-(90-78)*9 = $6990

When the demand is 79 and above the profit, while stacking 90 is at least 7000. Profit = 79*91-(90-79)*9 = $7090

Hence when the demand >=79 the profit is at least $7000.

The probability that the demand is >=79 is 0.5594 (use excel formula =1-POISSON.DIST(78,80,TRUE))

d. the quantity to at which profit is >$6000 is 69 = (69*91-(90-69)*9)

From above we know that at 78 the profit is less than $7000

The required probability is when Qty is between 69 and 78 which is 0.3436 [ use the formula =POISSON.DIST(78,80,TRUE) - POISSON.DIST(68,80,TRUE)]